# Verifying LLM Inference to Prevent Model Weight Exfiltration

### v1

@@ H.2 Context Length Ablations
To understand how input context length affects the GM-LS verification system’s performance, we conducted ablation studies using prompts of varying lengths. We generated prompts with 4,000, 8,000, and 12,000 tokens using the LongBench v2 dataset [BTZ+25], which provides realistic long-context tasks for multitask benchmarks. We create generations of max-length 500 tokens for 100 prompts, given 100 GM-LSthresholds, and $\sigma\in(0.001,0.01,0.1,1.)$. For each context length and model combination, we measured the trade-off between false positive rate and extractable information under the same experimental setup as Section 6.
@@
We draw this parallel as a potential future direction for developing future exfiltration mitigation. Current watermarking schemes are built for different infrastructure and generally designed to be robust to perturbations [KGW+23b, CGZ24, CHS25]. However, it is an interesting possibility for future work to adapt techniques from this literature to this setting.
@@ Appendix I Non-Exhaustive Taxonomy of Exfiltration Vectors
In this section we outline representative attack surfaces for exfiltrating model weights or related sensitive assets. For each surface, we describe how an attacker may operate and then give concrete scenarios with the attacker’s objective.
@@ I.1 Main Message Channel
An attacker may transmit secrets through the model’s ordinary user-facing outputs. In steganographic natural text, seemingly benign responses embed bits (e.g., via synonym choice or punctuation patterns) that the attacker later decodes.
@@ I.2 Developer Channels
An attacker may exploit development and maintenance environments—such as build machines, CI systems, and fine-tune hosts—to move model artifacts outside the secure boundary by abusing trusted workflows or misconfigurations.
  1. 1.
    Checkpoint copy to cloud. Scenario: A malicious or compromised developer transfers full checkpoints to external storage (e.g., S3) using valid credentials. Goal: Extract complete weights off-premises.
  2. 2.
    Insecure fine-tune / data leakage. Scenario: A fine-tune job on a shared cluster writes intermediates (weights, optimizer states, logs) to world-readable paths. Goal: Recover partial or full model states.
  3. 3.
    Insider exfiltration via tooling. Scenario: Developer tooling (debuggers, REPLs, laptops) is used to pull shards or artifacts via USB, SCP, or personal cloud. Goal: Smuggle model assets through routine workflows.
@@ I.3 Operational Security Channel
An attacker may target auxiliary systems used for backup, logging, and operational continuity to retrieve archived model artifacts. A representative case is backup key theft / restore abuse: the attacker compromises a backup service account or intercepts restore credentials and retrieves stored checkpoints. The objective is to decrypt or restore weights from offsite storage.
I.4 Side Channels (High-Level)
An attacker may exploit physical or low-level signals that correlate with internal computations to infer sensitive information; the literature documents numerous attacks and defenses. Examples include:
  1. 1.
    Timing channels. Scenario: Per-token or total latency variations correlate with secret-dependent execution paths. Goal: Infer bits from runtime patterns.
  2. 2.
    Power / electromagnetic channels. Scenario: Power draw or EM emissions reveal activity of model components. Goal: Reconstruct internal states or routing decisions.
  3. 3.
    Other ancillary signals. Scenario: Disk I/O patterns, cache contention, or network metadata are observed and correlated with secrets. Goal: Extract information from non-traditional observables.
@@ I.5 Miscellaneous
An attacker may combine internal services to bypass perimeter controls when the vector does not fit other categories. A representative example is Repeat-After-Me (internal cross-server exfiltration): an internal service with weight access sends “repeat after me: <binary payload>” to an externally exposed service, which echoes it to a client. The objective is to bridge a trusted internal boundary to an external sink without external prompting.
@@
    Monitor main traffic (this scheme). Score tokens for risk and send only the riskiest cases to extra checks; apply gentle throttles when needed. Examples: run a second, heavier check; re-generate an answer; temporarily pause streaming; slow or cap request rates.
@@
Figure 26: Egress-limiting schematic (illustrative): controls on outbound channels reduce bandwidth and increase observability for exfiltration attempts. Egress limiting as a form of strong guarantees on exfiltration is employed by the company Anthropic [Ant25].

### v2

@@ H.2 Context Length Ablations
To understand how input context length affects the FSSL-GM verification system’s performance, we conducted ablation studies using prompts of varying lengths. We generated prompts with 4,000, 8,000, and 12,000 tokens using the LongBench v2 dataset [bai2025longbenchv2deeperunderstanding], which provides realistic long-context tasks for multitask benchmarks. We create generations of max-length 500 tokens for 100 prompts, given 100 FSSL-GMthresholds, and $\sigma\in(0.001,0.01,0.1,1.)$. For each context length and model combination, we measured the trade-off between false positive rate and extractable information under the same experimental setup as Section 6.
@@
We draw this parallel as a potential future direction for developing future exfiltration mitigation. Current watermarking schemes are built for different infrastructure and generally designed to be robust to perturbations [ARXIV:KGWKMG23, COLT:ChrGunZam24, SP:CHS25]. However, it is an interesting possibility for future work to adapt techniques from this literature to this setting.
@@ Appendix I Non-Exhaustive Taxonomy of Exfiltration Vectors
In this section we outline a handful of attack surfaces for exfiltrating model weights (or related sensitive data). We describe how an attacker may operate and give concrete example.
@@ I.1 Main Message Channel
An attacker may transmit secrets through the model’s ordinary user-facing outputs. In steganographic natural text, seemingly benign responses embed bits (e.g., via synonym choice or punctuation patterns) that the attacker later decodes. This is the primary focus of this work.
@@ I.2 Developer Channels
An attacker may exploit development or maintenance environments—such as build machines, CI systems, and fine-tune hosts—to exfiltrate model weights or sensitive data by abusing trusted workflows or misconfigurations.
This could happen if a a malicious or compromised developer can transfer full checkpoints to external storage (e.g., AWS S3) using valid credentials. Or less maliciously, if a finetuning job on a shared cluster can write intermediates (weights, optimizer states, logs) to world-readable paths.
@@ I.3 Operational Security Channel
An attacker may target systems used for backup, logging, and operational continuity to retrieve archived model artifacts. For example, the attacker could compromise a backup service and make a copy of the weights, which would likely be encrypted, from there they only need to exfiltrate a small key to decrypt the backups from the main development server. As such, it is critical to protect not just the datacenter serving weights, but also servers storing auxiliary information, like backups, or key management.
I.4 Side Channels
An attacker outside of a datacenter may exploit side-channels that correlate with internal computations to infer sensitive information; or an attacker inside of a datacenter may send information outside encoded through a “side-channel”. For example, for timing channels: per-token or total latency variations correlate with secret-dependent execution paths, or timing delays could be used by a adversary to send a specifically encoded message.
@@ I.5 Miscellaneous
There are also a suite of attacks that are either LLM specific or unique to the workflow associated with an AI workflow. Notably, defending against a distillation attack, where an adversary outside of the datacenter trains their model on the outputs of a closed source model, requires an entirely different set of defenses.
Additionally, an attacker may combine internal services to bypass perimeter controls when the vector does not fit other categories. An example is Repeat-After-Me (internal cross-server exfiltration) attack: an internal service with weight access sends “repeat after me: <binary payload>” to an externally exposed service, which echoes it to a client. In such a setting, the 2nd server is doing expected behavior and would pass verification. In spirit, this is akin to a confused deputy problem.
@@
    Monitor main traffic (this scheme). Score tokens for risk and send only the riskiest cases to extra checks; likely apply rate-limiting when needed.
@@
Figure 26: Egress-limiting schematic (illustrative): controls on outbound channels reduce bandwidth and increase observability for exfiltration attempts. Egress limiting as a form of strong guarantees on exfiltration is employed by the company Anthropic [anthropic2025asl3].